Ganymede's orbit around Jupiter is nearly circular and has a period of 7.155 days. Assume Ganymede is moving in uniform circular motion.
Answer #1

what I have to find radius or angular speed or angular acceleration???

from the given data U did'nt mention what I have to find so Am finding all the three...

The period is the time for Ganymede's to make one orbit around Jupiter. Let’s convert it to seconds.

t = 7.155 days * 24 hr/day * 60 min/hr * 60 s/min = 6.18*105 seconds

We can use the following equation to determine the radius of the orbit.
T2 = [4π2 / (G*M)]*r3

Mass of Saturn = 1.898*1027 kg
G * M = 6.67*10-11 * 1.898*1027 = 1.26*1017
2 / (G*M)] = 3.11*10-16
r^3 = (6.18*105)2 / 3.11*10-16
r^3 = 1.22*1027 m3 =>

r = 1.07*109 m


Orbital speed can be calculated as follows:

V₀ = √(GM/r) = sqrt[1.26*1017 / 1.07*109] = 1.08*104 m/s

Centripetal acceleration = v2/r = 0.110 m/s2